OpenStudy (anonymous):

Given the function f(x) = 5(x+4) - 6, solve for the inverse function when x = 19. a. 1 b. 68 c. 72 d. 84

OpenStudy (anonymous):

solve \[5(x+4) - 6=19\]

OpenStudy (anonymous):

5x+20-6=19 5x+14=19 5x=5 x=1???

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

twas me, i would have written \[5(x+4) - 6=19\\ 5(x+4)=25\\ x+4=5\\x=1\] but same answer